OpenStudy (anonymous):

PLEASE HELP MEDAL GRANTED which ordered pair is the vertex of y= x+5 -1 A (–5, –1) B(–1, –5 ) C(5, –1) D(–1, 5) which ordered pair is the vertex of y= x-3 +2 A (3, 2) B (2, –3) C (–3, 2) D (2, 3)

OpenStudy (campbell_st):

can I just check... are the 2 equations \[y = (x +5)^2 -1\] and \[y = (x -3)^2 + 2\]